Question 3 (1 point) 3 An object is removed from a room where the temperature is 71 degrees and is taken outside, where the air temperature is 29 dėgrees. After 1 minute, the temperature of the object reads 53 degrees. What will be the temperature of the object att 2 minutes? (round your answer to two decimal places) dT HINT: Use the equation di-k (7-7,) Your Answer Answer Question 4 (1 point) The population in a town of 60,000 is growing by 9% per year. How long will it take the population to grow to 190,000? Round to the nearest year. Your Answer: Answer
Answer #1

Hey,

Note: Brother in case of any queries, just comment in box I would be very happy to assist all your queries

1)

dT/(T-Ta)=-kdt

log(53-29)-log(71-29)=-k

k=log(1.75)

So, for t=22

dT/(T-Ta)=-kdt

log(T-29)-log(71-29)=-2k

So, T=42.714285714285715 degrees

2)

((1.09)^t)*60000

So, t=14 years

Kindly revert for any queries

Thanks.